You are asked to value Sunshine Inc. using the relative valuation method. Sunshine Inc.'s earnings forecast for next year (EPS (next year)) is $2.44. The valuation and earnings of comparable companies are provided below. What is your estimate for the company's stock price? Round to one decimal place. ​[Hint: Compute the corresponding valuation multiple of the comparables and take simple average across the three comparable companies. Then apply this average multiple to Sunshine Inc.] answerable question reference

answerable question reference
Answer #1

average multiple=((18.3/1.22)+(26.3/1.42)+(55.9/2.80))/3=17.82845

your estimate for the company's stock price=2.44*17.8285=43.50

the above will be answer..